OpenStudy (anonymous):

given that cos theta=4/5find sin theta

OpenStudy (dumbcow):

use trig identity sin^2+cos^2 =1 sin^2 = 1-cos^2 ....... since cos = 4/5 cos^2 = 16/25 1-cos^2 = 1-16/25 = 9/25 ......... sin^2 = 9/25 sin = 3/5

OpenStudy (anonymous):

Or you could use a right triangle. Knowing cosine = adj/hyp, you could use pythagorean theorem (a^2 + b^2 = c^2) to find the missing side of the triangle (which would be 3). Then since, sine = opposite/hyp, you would get 3/5